# Nonparametric Estimation of Joint Entropy through Partitioned Sample-Spacing Method

### v1

@@ 4.1 Multivariate Density Estimator
$\frac{n_{k}}{n}\cdot\prod_{i=1}^{d}\frac{2m_{k}}{n_{k}(x_{i,(a_{i}+m_{k})}^{k}-x_{i,(a_{i}-m_{k})}^{k})}$
@@
Proposition 3.
The partitioned multivariate sample-spacing density estimator $\hat{f}_{n,l}(\mathbf{x})$ satisfies
$\lim_{n\rightarrow\infty}\int_{\mathbb{R}^{d}}\hat{f}_{n,l}(\mathbf{x})\,d\mathbf{x}=1,$
Now that we have established that $\hat{f}_{n,l}(\mathbf{x})$ is an asymptotically valid probability density function, we proceed to the main result for the estimator: proving that it is also strongly consistent. This means showing that the estimator converges almost surely to the true underlying density $f(\mathbf{x})$ as the sample size increases. The following theorem formalizes this crucial property.
Theorem 4 (Consistency of the Partitioned Multivariate Density Estimator).
@@
$\hat{f}_{n,\ell}(\mathbf{x})=\frac{n_{k}}{n}\cdot\prod_{j=1}^{d}\hat{f}_{X_{j};k}(x_{j})\text{}$
where $\hat{f}_{X_{j};k}(x_{j})=\frac{2m_{k}}{n_{k}(x_{j,(a_{j}+m_{k})}^{k}-x_{j,(a_{j}-m_{k})}^{k})}$. Assume that as $n\to\infty$, the number of partitions satisfies $\ell(n)\to\infty$ with $\ell(n)^{d}=o(n)$. Further, assume the spacing parameter $m_{k}$ satisfies the conditions $m_{k}\to\infty$ and $m_{k}/n_{k}\to 0$, which are met by the globally adopted rate of $m_{k}=\lfloor\sqrt{n_{k}}+1/2\rfloor$. Given these conditions, $\hat{f}_{n,\ell}(\mathbf{x})$ converges almost surely to the true joint density $f(\mathbf{x})$:
$\hat{f}_{n,\ell}(\mathbf{x})\xrightarrow{a.s.}f(\mathbf{x})$
@@ 4.2 Multivariate Joint Entropy Estimator
$H_{v}(\mathbf{X},n,\ell)=-\frac{1}{n}\sum_{v=1}^{n}\log\hat{f}_{n,\ell}(\mathbf{X}_{v}).$
@@
$\log\hat{f}_{n,\ell}(\mathbf{X}_{v})=-\log\left(\frac{n_{k}}{n}\cdot\prod_{j=1}^{d}\frac{2m_{k}}{n_{k}\Delta x_{j,a_{j}}^{k}}\right)$
@@
Theorem 5 (Consistency of the Entropy Estimator $\hat{H}_{n,\ell}$).
Assume that the true density $f$ satisfies $\log f\in L^{1+\delta}$ for some $\delta\in(0,1)$, $m_{k}=\lfloor\sqrt{n_{k}}+1/2\rfloor$, with $n_{k}\geq 1$, and that the number of partitions satisfies $\ell(n)^{d}=o(n)$. Then, the entropy estimator $\hat{H}_{n,\ell}$ converges in $L^{1}$ to the true entropy $H(f):=\mathbb{E}[-\log f(X)]$ as $n\to\infty$:
@@
Building upon the $L^{1}$ and almost sure convergence of the PSS entropy estimator established in Theorem 5, we can immediately obtain the consistency of the corresponding mutual information estimator.
Corollary 2 (Consistency of the PSS Mutual Information Estimator).
@@
$\hat{I}_{n,\ell}(X;Y)\xrightarrow{L^{1}}I(X;Y),\ \hat{I}_{n,\ell}(X;Y)\xrightarrow{a.s.}I(X;Y)$
@@ 5 Numerical Experiments
We compare PSS against kkNN estimators (KL, KSG), CADEE, and uniformizing-mapping versions (UM-tKL, UM-tKSG). To ensure a fair comparison, the key hyperparameter for each method—the partition parameter $\ell$ for PSS and the number of neighbors $k$ for kNN estimators—was chosen to minimize the empirical Mean Squared Error (MSE) in each setting. All results are averaged over 100 trials and reported as rooted mean squared error (RMSE) relative to the true entropy. The standard kkNN estimator has a naive time complexity of $O(N^{2}d)$ in $d$ dimensions, since all pairwise distances are computed. Tree-based methods (k-d tree, Ball tree) reduce query cost to $O(Nd\log N)$ after an $O(Nd\log N)$ build phase, but their efficiency deteriorates with dimension and skewness. Normalizing flow (NF) estimators AoLi2022 require training a deep generative model with $O(Nd)$ forward/backward passes per epoch, repeated over many epochs, leading to substantial computational overhead. In contrast, PSS assigns $N$ samples to $\ell^{d}$ partitions in $O(Nd)$, sorts within partitions in about $O(dN\log(N/\ell^{d}))$, and computes entropy in $O(Nd)$. As $\ell$ grows with $N$, the factor $N/\ell^{d}$ shrinks and the cost approaches $O(Nd)$, yielding favorable scalability. CADEE copula runs in reasonable time for small problems but its runtime grows rapidly with NN and $d$ relative to PSS, and its accuracy degrades in higher dimensions.
@@
We first consider multivariate normal data with $\rho=0$. Figure 2 reports RMSE and runtime across varying sample sizes ($d=10$) and dimensions ($N=3000$). PSS achieves consistently lower RMSE than kkNN and CADEE, with the advantage widening as $d$ grows. In terms of runtime, PSS scales nearly linearly with NN and remains substantially faster than CADEE, while also outperforming kkNN methods. These results confirm that PSS mitigates the curse of dimensionality in Gaussian settings while maintaining computational efficiency.
Figure 3: RMSE (a, c) and runtime (b, d) for entropy estimators under the Gamma distribution with shape =0.4=0.4, scale =0.3=0.3. Panels (a, b): $d=5$, varying sample size NN. Panels (c, d): $\rho=0$ $N=30000$, varying dimension.
@@
where $h(X_{j})$ is the marginal Gamma entropy and $R$ is the copula correlation matrix. PSS remains highly accurate under skewed and heavy-tailed marginals, substantially outperforming kkNN and CADEE at moderate $d$. Runtime comparisons again highlight the scalability of PSS relative to CADEE, whose cost grows steeply with NN and $d$. These findings demonstrate that PSS adapts effectively to skewness and bounded supports while preserving its computational advantage.
6 Discussion
We introduced the partitioned sample-spacing (PSS) estimator, a nonparametric method for joint entropy. Theoretical guarantees include almost sure and $L^{1}$ consistency. Empirically, PSS outperforms kkNN and related methods in moderate to high dimensions and adapts well to skewed distributions, while scaling efficiently with sample size. Its efficiency stems from local partitioning, which avoids exhaustive pairwise distance computations. The partition parameter $\ell$ is central: larger $\ell$ improves accuracy at modest additional cost, though automated selection remains an open problem. As with other grid-based methods, the exponential growth of partitions limits applicability in very high dimensions. Indeed, as seen in Figures 2c and 3c, performance can degrade for larger values of l in very high dimensions $(d>20)$, suggesting that overly fine partitions may lead to unstable local estimates in sparse settings.
A natural avenue for addressing the choice of $\ell$ is cross-validation. One may split the data into training and validation folds, fit $\hat{f}_{n,\ell}$ on the training fold, and evaluate the average negative log-density on the validation fold. Since our entropy estimator itself is defined as an empirical average of $\log\hat{f}_{n,\ell}(X)$, this validation criterion is directly aligned with the estimation objective. In effect, CV minimizes a finite-sample surrogate of the Kullback–Leibler risk, suggesting that the selected $\ell$ may asymptotically approach the optimal rate. Compared to kkNN-based estimators, which often require ad hoc bias–variance heuristics, and NF-based estimators, which demand expensive model training, PSS admits a simple and principled likelihood-based validation strategy. Developing formal guarantees for such data-driven selection remains an important future direction.
Overall, PSS offers a consistent, scalable alternative to kkNN. Future work includes characterizing its bias–variance trade-off and developing more principled strategies for tuning $\ell$, which would enhance its utility in machine learning and information-theoretic applications.
7 Additional Experiment: Evaluating Independent Component Analysis on Real-World EEG Data via Cross-Validation
@@ 6 Discussion
We used the UCI EEG Eye State dataset, a 14-dimensional time-series signal comprising N=14,980N=14{,}980 samples and $d=14$ channels. The dataset is a continuous recording from a single subject. To prepare the data for entropy estimation—which assumes independent and identically distributed (i.i.d.) samples—we first pre-processed the entire time series via whitening, which zero-centers the data and applies a linear transformation to decorrelate the features. Each row of the transformed data matrix was then treated as an i.i.d. sample from a stationary $d$-dimensional distribution. This simplification ignores temporal dependence and is adopted solely for evaluating entropy estimators on real-world correlated data.
@@
We applied the FastICA algorithm to the whitened data ($\mathbf{X}_{w}$) to separate it into a set of statistically independent components ($\mathbf{Y}$). All computations were performed using scikit-learn 1.5.2 (FastICA, whiten=True) on an Apple M2, macOS 14.6, Python 3.9 environment.
@@
The results are visualized in Figure 4 and summarized in Table 1. These highlight the performance of PSS relative to several widely used $k$-NN based estimators, including KL, KSG, and their truncated variants (tKL, tKSG).
@@
8 Proofs
8.1 Proof of lemma 1

### v2

@@ 4.1 Multivariate Density Estimator
$\hat{f}_{n,\ell}(\mathbf{x})=\frac{n_{k}}{n}\cdot\prod_{i=1}^{d}\frac{2m_{k}}{n_{k}(x_{i,(a_{i}+m_{k})}^{k}-x_{i,(a_{i}-m_{k})}^{k})}$
@@
Remark.
The local spacing-based density is defined only for partitions with $n_{k}\geq 2$, since valid $m$-spacings require at least two sample points in the cell. Partitions with $n_{k}=0$ or $n_{k}=1$ are skipped and do not contribute to $\hat{f}_{n,\ell}(X_{v})$ or to the entropy estimator.
Proposition 9
The partitioned multivariate sample-spacing density estimator $\hat{f}_{n,\ell}(\mathbf{x})$ satisfies
$\lim_{n\rightarrow\infty}\int_{\mathbb{R}^{d}}\hat{f}_{n,\ell}(\mathbf{x})\,d\mathbf{x}=1.$
See Appendix H for the proof sketch. Now that we have established that $\hat{f}_{n,\ell}(\mathbf{x})$ is an asymptotically valid probability density function, we proceed to the main result for the estimator: proving that it is also strongly consistent. This means showing that the estimator converges almost surely to the true underlying density $f(\mathbf{x})$ as the sample size increases. The following theorem formalizes this crucial property.
Theorem 10 (Consistency of the Partitioned Multivariate Density Estimator)
@@
$\hat{f}_{n,\ell}(\mathbf{x})=\frac{n_{k}}{n}\cdot\prod_{j=1}^{d}\widehat{f}_{X_{j};k}(x_{j}),$
where $\widehat{f}_{X_{j};k}(x_{j})=\frac{2m_{k}}{n_{k}(x_{j,(a_{j}+m_{k})}^{k}-x_{j,(a_{j}-m_{k})}^{k})}$. Assume that as $n\to\infty$, the number of partitions satisfies $\ell(n)\to\infty$ with $\ell(n)^{d}=o(n)$. Further, assume the spacing parameter $m_{k}$ satisfies the conditions $m_{k}\to\infty$ and $m_{k}/n_{k}\to 0$, which are met by the globally adopted rate of $m_{k}=\lfloor\sqrt{n_{k}}+1/2\rfloor$. Given these conditions, $\hat{f}_{n,\ell}(\mathbf{x})$ converges almost surely and in $L^{1}$ to the true joint density $f(\mathbf{x})$:
$\hat{f}_{n,\ell}(\mathbf{x})\xrightarrow{a.s.}f(\mathbf{x})\quad\text{and}\quad\hat{f}_{n,\ell}(\mathbf{x})\xrightarrow{L^{1}}f(\mathbf{x}).$
See Appendix I for the proof sketch.
@@ 4.2 Multivariate Joint Entropy Estimator
$\hat{H}_{n,\ell}(\mathbf{X})=-\frac{1}{n}\sum_{v=1}^{n}\log\hat{f}_{n,\ell}(\mathbf{X}_{v}).$
@@
$\log\hat{f}_{n,\ell}(\mathbf{X}_{v})=-\log\left(\frac{n_{k}}{n}\cdot\prod_{j=1}^{d}\frac{2m_{k}}{n_{k}\Delta x_{j,a_{j}}^{k}}\right),$
@@
Theorem 11 (Consistency of the Entropy Estimator $\hat{H}_{n,\ell}$)
Assume that the true density $f$ satisfies $\log f\in L^{1+\delta}$ for some $\delta\in(0,1)$, $m_{k}=\lfloor\sqrt{n_{k}}+1/2\rfloor$, with $n_{k}>1$, and that the number of partitions satisfies $\ell(n)^{d}=o(n)$. Then, the entropy estimator $\hat{H}_{n,\ell}$ converges in $L^{1}$ to the true entropy $H(f):=\mathbb{E}[-\log f(X)]$ as $n\to\infty$:
@@
See Appendix J for the proof sketch. Building upon the $L^{1}$ and almost sure convergence of the PSS entropy estimator established in Theorem 11, we can immediately obtain the consistency of the corresponding mutual information estimator.
Corollary 12 (Consistency of the PSS Mutual Information Estimator)
@@
$\hat{I}_{n,\ell}(X;Y)\xrightarrow{L^{1}}I(X;Y),\quad\hat{I}_{n,\ell}(X;Y)\xrightarrow{a.s.}I(X;Y).$
4.3 Data-Driven Parameter Selection via Cross-Validation
A critical challenge in nonparametric estimation is the selection of the smoothing parameter—in our case, the number of partitions per dimension, $\ell$. While theoretical analyses for grid-based estimators suggest that the partition size should scale with the sample size, the optimal rate typically depends on unknown properties of the true density $f$ (e.g., roughness or derivatives). To address this in a practical, data-driven manner without relying on unverifiable assumptions, we propose a Likelihood-based Cross-Validation strategy.
Let the dataset $\mathcal{D}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}$ be partitioned into $K$ disjoint folds. For each fold $k\in\{1,\dots,K\}$, let $\mathcal{V}_{k}$ denote the validation set and $\mathcal{T}_{k}=\mathcal{D}\setminus\mathcal{V}_{k}$ the training set. We select the optimal parameter $\ell^{*}$ by minimizing the average negative log-likelihood on the hold-out data:
$\ell^{*}=\operatorname*{argmin}_{\ell}\left\{-\frac{1}{K}\sum_{k=1}^{K}\frac{1}{|\mathcal{V}_{k}|}\sum_{\mathbf{x}\in\mathcal{V}_{k}}\log\hat{f}_{n,\ell}^{(\mathcal{T}_{k})}(\mathbf{x})\right\},$ (1)
where $\hat{f}_{n,\ell}^{(\mathcal{T}_{k})}$ represents the PSS density estimator constructed using only the training subset $\mathcal{T}_{k}$.
Connection to the Entropy Objective.
The proposed cross-validation criterion is not merely a heuristic but is intrinsically linked to the objective of entropy estimation. Recall that the plug-in entropy estimator is defined as the empirical average of the negative log-density:
$\hat{H}_{n,\ell}=-\frac{1}{n}\sum_{i=1}^{n}\log\hat{f}_{n,\ell}(\mathbf{x}_{i})\,.$ (2)
Comparing Eq. (1) with the definition of $\hat{H}_{n,\ell}$, it is evident that the validation loss $\mathcal{L}_{CV}$ takes an identical form to the entropy estimator itself. By the law of large numbers, the expectation of this loss converges to the cross-entropy between the true density $f$ and the estimator $\hat{f}$:
$\mathbb{E}_{\mathbf{x}\sim f}\left[-\log\hat{f}_{n,\ell}(\mathbf{x})\right]=H(f)+D_{KL}(f\,||\,\hat{f}_{n,\ell})\,.$ (3)
Since the true entropy $H(f)$ is constant with respect to $\ell$, minimizing the CV loss is asymptotically equivalent to minimizing the Kullback-Leibler (KL) divergence $D_{KL}(f\,||\,\hat{f}_{n,\ell})$. This structural alignment ensures that the hyperparameter $\ell$ is optimized specifically for the goal of accurate information estimation, providing a coherent and principled framework for automated model selection.
4.4 Practical Implementation Note
The asymptotic condition $\ell(n)^{d}=o(n)$ in our theoretical analysis is a sufficient but not a necessary requirement for consistency. In practical applications, the PSS estimator evaluates $\hat{f}_{n,\ell}(X_{v})$ only at observed sample points; partitions with $n_{k}=0$ or with insufficient spacing do not contribute to the entropy computation. As a result, the estimator is governed by the number of occupied partitions rather than by the full $\ell^{d}$ grid.
Real datasets, including the whitened 14-dimensional EEG signals used in our experiments, typically concentrate on a lower-dimensional manifold. Only a small fraction of the $\ell^{d}$ hyperrectangles actually contain data, and the entropy estimator acts only on these populated regions. Consequently, practical choices such as $\ell^{\ast}=12$ remain stable and do not conflict with the asymptotic condition, since the estimator adapts to the effective geometry and occupancy pattern of the data.
Algorithm 1 PSS Joint Entropy Estimation
1:Samples $S=\{\mathbf{x}_{v}\}_{v=1}^{n}\subset\mathbb{R}^{d}$, partition parameter $\ell$.
2:Estimated joint entropy $\hat{H}_{n,\ell}$.
3:Partition each axis into $\ell$ equal-width intervals to form $\ell^{d}$ cells $\{P_{k}\}$.
[…]
6:for each occupied partition $P_{k}$ with $n_{k}\geq 2$ do
7:  Sort points in $P_{k}$ marginally.
8:  Compute spacings $\Delta x^{k}_{j,a_{j}}=x^{k}_{j,(a_{j}+m_{k})}-x^{k}_{j,(a_{j}-m_{k})}$ using $m_{k}=\lfloor\sqrt{n_{k}}\rfloor$.
9:end for
[…]
11:for $v=1$ to $n$ do
12:  Identify partition $P_{k}$ containing $\mathbf{x}_{v}$.
13:  if $n_{k}<2$ or any $\Delta x^{k}_{j,a_{j}}=0$ then
14:   continue $\triangleright$ Skip points with undefined density
15:  end if
16:  Calculate local log-density contribution:
$L_{v}=\log\left(\frac{n_{k}}{n}\right)+\sum_{j=1}^{d}\log\left(\frac{2m_{k}}{n_{k}\,\Delta x^{k}_{j,a_{j}}}\right)$
17:  $\mathcal{S}\leftarrow\mathcal{S}+L_{v}$
18:end for
[…]
20:return $\hat{H}_{n,\ell}$
@@ 5 Numerical Experiments
To comprehensively evaluate the performance and practical utility of the proposed Partitioned Sample-Spacing (PSS) estimator, we conducted extensive experiments on both synthetic and real-world datasets. Our experimental design focuses on validating three key properties: (1) robustness to skewed distributions, (2) computational scalability in large-sample regimes, and (3) reliability in high-dimensional downstream tasks.
Hyperparameter Selection.
For hyperparameter selection—specifically, the partition parameter $\ell$ for PSS and the number of neighbors $k$ for kNN—we adopted distinct protocols depending on the availability of ground truth:
  • Synthetic Benchmarks (Oracle Tuning): In scenarios where the true probability density and entropy $H(f)$ are known, we selected hyperparameters that minimized the MSE. This allows us to assess the empirical bound of the estimator’s accuracy and convergence properties, independent of hyperparameter selection error.
  • Real-World Applications (Data-Driven Tuning): For real-world tasks such as ICA and feature selection, where the ground truth is inaccessible, we employed the Likelihood-based Cross-Validation(CV) strategy proposed in Section 4.3. This evaluates the estimators in a realistic setting, demonstrating their stability and reliability when fully automated, data-driven tuning is required.
5.1 Synthetic Data Experiments
We compare PSS against kkNN estimators (KL, KSG), copula based CADEE (copula), and uniformizing-mapping versions (UM-tKL, UM-tKSG) AoLi2022. For both CADEE and the UM-based estimators, we adapted the authors’ official implementations to fit our experimental framework. We do not include general NF-based entropy estimators as baselines. As shown in (AoLi2022), pure normalizing-flow estimators are typically less accurate and less stable than the normalized truncated kkNN estimators. To ensure a fair comparison, the key hyperparameter for each method—the partition parameter $\ell$ for PSS and the number of neighbors $k$ for kNN estimators—was chosen to minimize the empirical Mean Squared Error (MSE) in each setting. All results are averaged over 100 trials and reported as rooted mean squared error (RMSE) relative to the true entropy.
Computational Complexity Analysis.
The standard kkNN estimator has a naive time complexity of $O(N^{2}d)$ in $d$ dimensions, since all pairwise distances are computed. Tree-based methods (k-d tree, Ball tree) reduce query cost to $O(Nd\log N)$ after an $O(Nd\log N)$ build phase, but their efficiency deteriorates with dimension and skewness. Normalizing flow (NF) estimators (AoLi2022) require training a deep generative model with $O(Nd)$ forward/backward passes per epoch, repeated over many epochs, leading to substantial computational overhead. In contrast, PSS assigns $N$ samples to $\ell^{d}$ partitions in $O(Nd)$, sorts within partitions in about $O(dN\log(N/\ell^{d}))$, and computes entropy in $O(Nd)$. As $\ell$ grows with $N$, the factor $N/\ell^{d}$ shrinks and the cost approaches $O(Nd)$, yielding favorable scalability. CADEE (copula) runs in reasonable time for small problems but its runtime grows rapidly with NN and $d$ relative to PSS, and its accuracy degrades in higher dimensions.
Figure 3: RMSE (a, c) and runtime (b, d) for entropy estimators under the Gamma distribution with shape = 0.4, scale = 0.3. Panels (a, b): $d=5$, varying sample size NN. Panels (c, d): $\rho=0$, $N=30,000$, varying dimension.
Figure 4: RMSE (a, c) and runtime (b, d) for entropy estimators with varying correlation coefficient $\rho$. Panels (a, b): Normal distribution ($d=5,N=20,000$). Panels (c, d): Gamma distribution (shape =0.4=0.4, scale =0.3=0.3) with $d=7,N=50,000$.
@@ 5.1 Synthetic Data Experiments
We first consider multivariate normal data with $\rho=0$. Figure 2 reports RMSE and runtime across varying sample sizes ($d=10$) and dimensions ($N=3000$). In terms of accuracy (Figure 2a, c), PSS consistently outperforms traditional baselines (KL, KSG) and CADEE, with the performance gap widening as $d$ grows. Notably, PSS achieves accuracy competitive with state-of-the-art normalizing flow-based estimators (UM-tKL, UM-tKSG), effectively matching their performance even in higher dimensions ($d=40$) without the need for auxiliary density modeling. In terms of runtime (Figure 2b, d), PSS scales nearly linearly with NN and remains substantially faster than CADEE and the UM-variants, indicating that PSS alleviates several dimensionality-related challenges in Gaussian settings while maintaining strong computational efficiency.
@@
where $h(X_{j})$ is the marginal Gamma entropy and $R$ is the copula correlation matrix. In terms of accuracy, PSS exhibits remarkable robustness to skewed distributions and semi bounded supports. It significantly outperforms traditional baselines (KL, KSG) and CADEE, which suffer from severe boundary bias in this setting. Crucially, PSS matches or exceeds the performance of normalizing flow-based methods (UM-tKL, UM-tKSG), demonstrating that the partitioning strategy effectively captures local density structures without relying on learned transformations. Runtime comparisons further highlight the advantage of PSS; it scales linearly with NN and remains substantially faster than CADEE, while avoiding the computational overhead associated with training flow-based models.
Robustness under Correlation.
We further investigate the estimator’s sensitivity to dependence strength by varying the correlation coefficient $\rho\in[0,0.8]$ (Figure 4). In terms of accuracy (Figure 4a, c), PSS exhibits remarkable stability, maintaining low errors even as traditional baselines degrade rapidly with increasing correlation. Notably, in the high-correlation Gamma setting ($\rho=0.8$), PSS surpasses even the normalizing flow-based variants without requiring learned transformations. This robustness is directly attributable to the adaptive behavior of the optimal partition parameter $\ell^{*}$. We observe that a larger $\ell^{*}$ is required to minimize error as correlation strengthens (e.g., increasing from 4 to 21 in the Gamma case), indicating that a finer partitioning is essential to capture the probability mass concentrating in narrow regions. Regarding computational cost (Figure 4b, d), PSS displays a flat runtime profile, confirming that its complexity remains invariant to the dependence structure, whereas CADEE suffers significant slowdowns. However, we note a structural limitation as $\rho\to 1$; while PSS handles $\rho=0.8$ robustly, the transition to a fully degenerate distribution on a lower-dimensional manifold remains a fundamental challenge for full-dimensional density estimators.
5.2 Real-World Data Experiments
5.2.1 ICA Evaluation Experiment
@@
We used the UCI EEG Eye State dataset(eeg_eye_state_264), a 14-dimensional time-series signal comprising N=14,980N=14{,}980 samples and $d=14$ channels. The dataset is a continuous recording from a single subject. To prepare the data for entropy estimation—which assumes independent and identically distributed (i.i.d.) samples—we first pre-processed the entire time series via whitening, which zero-centers the data and applies a linear transformation to decorrelate the features. Each row of the transformed data matrix was then treated as an i.i.d. sample from a stationary $d$-dimensional distribution. This simplification ignores temporal dependence and is adopted solely for evaluating entropy estimators on real-world correlated data.
@@
We applied the FastICA algorithm to the pre-whitened data ($\mathbf{X}_{w}$) to separate it into a set of statistically independent components ($\mathbf{Y}$). All computations were performed using scikit-learn 1.5.2 (FastICA) on an Apple M3 processor.
@@
The results are visualized in Figure 5 and summarized in Table 1. These highlight the performance of PSS relative to several widely used $k$-NN based estimators, including KL, KSG, and their truncated variants using normalizing flow (tKL, tKSG).
@@
5.2.2 Application to Feature Selection
In this experiment, we evaluate the practical utility of PSS for feature selection in high-dimensional data. Unlike static entropy estimation, feature selection requires iteratively evaluating mutual information for hundreds of candidate subsets, demanding both computational efficiency and ranking stability. Neural MI estimators, while powerful, require repeated neural-network optimization and are therefore computationally expensive in feature-selection settings (Sreejith), so we restrict our comparison to nonparametric estimators.
1. Dataset Description.
We utilized the UCI Appliances Energy Prediction dataset
(appliances_energy_prediction_374), a widely used benchmark for high-dimensional regression and classification. The dataset consists of N=19,735N=19{,}735 samples and $d=26$ continuous features, including sensors for temperature (T1–T9), humidity (RH_1–RH_9), and weather conditions. The features exhibit strong multicollinearity (e.g., correlations between indoor sensors), posing a challenge for distance-based estimators. To frame this as a classification task for mutual information estimation $I(\mathbf{S};Y)$, we discretized the target variable (energy consumption) into binary classes (High/Low) based on its median.
2. Experimental Setup.
We performed Greedy Forward Selection to rank the features. Starting with an empty set, at each step $t$, we selected the feature $X_{j}$ that maximized the estimated mutual information with the target $Y$, conditioned on the set $\mathbf{S}_{t-1}$:
$X^{*}=\operatorname*{argmax}_{X_{j}\notin\mathbf{S}_{t-1}}\hat{I}(\mathbf{S}_{t-1}\cup\{X_{j}\};Y)\,.$
We compared PSS against the $k$-NN based KL estimator. For PSS, the optimal partition parameter selected via cross-validation was $\ell^{*}=2$, which effectively adapts to the sparsity of the high-dimensional joint space. Note that neural estimation methods (e.g., MINE, Normalizing Flows) were excluded from this experiment due to their prohibitive computational cost and stochastic instability in iterative search processes. For evaluation, the selected feature subsets were fed into a Support Vector Machine (SVM) classifier with a Radial Basis Function (RBF) kernel. The classification accuracy was measured on a held-out test set (30%).
3. Results and Analysis.
Figure 6 presents the results. Figure 6(a) shows the classification accuracy as features are added. PSS (blue) consistently identifies more informative features in the early to middle stages compared to kNN (red), achieving higher accuracy with fewer variables.
More importantly, Figure 6(b) reveals the theoretical behavior of the estimators. As the dimensionality of the selected subset increases, the estimated MI should fundamentally be non-decreasing (Data Processing Inequality). PSS adheres to this property, showing a stable monotonic increase. In contrast, the kNN estimator collapses after approximately 5 features, producing decreasing and even negative MI values. This confirms that distance-based estimators suffer from severe bias in high-dimensional subspaces ($d>5$) due to the scarcity of samples, whereas the partitioning strategy of PSS maintains robustness and theoretical consistency.
(a) Classification Accuracy (SVM)
(b) Estimated MI Stability
Figure 6: Feature selection results on the UCI Energy dataset ($d=26$). (a) SVM classification accuracy. PSS (blue) yields higher accuracy than kNN (red). (b) Estimated Mutual Information. Unlike kNN, which degrades into negative values, PSS maintains theoretical consistency (monotonic increase).
6 Discussion
We introduced the Partitioned Sample-Spacing (PSS) estimator, a nonparametric method for joint entropy estimation. By applying univariate spacing methods locally within partitions, PSS achieves almost sure and $L^{1}$ consistency while avoiding the $O(N^{2})$ distance computations of kNN methods. The resulting $O(Nd)$ complexity makes PSS competitive with neural estimators in runtime, without requiring training.
Our experiments show PSS performs well across diverse settings. In synthetic benchmarks on both Gaussian and skewed distributions, PSS consistently achieves lower MSE than kNN methods while matching the accuracy of state-of-the-art normalizing flow variants (UM-tKL, UM-tKSG). Notably, PSS demonstrates remarkable robustness under strong correlation, where its adaptive partitioning strategy effectively captures the concentrated probability mass that typically degrades the performance of global distance-based estimators. We used oracle tuning in these controlled experiments to establish PSS’s theoretical potential and enable fair comparison with optimally-tuned baselines.
Real-world applications reveal a critical advantage: PSS maintains theoretical validity where kNN estimators systematically fail. This aligns with the observation of (brown12a) that accurate MI estimation becomes unreliable beyond moderate dimensions, whereas PSS remains stable in exactly these regimes. Using cross-validation for hyperparameter selection, PSS produced non-negative total correlation values in ICA on EEG data ($d=14$), while kNN yielded physically impossible negative values—a fundamental violation of information theory. In feature selection on energy prediction data ($d=26$), PSS preserved the monotonic increase of mutual information as features were added, consistent with the data processing inequality, while kNN exhibited decreasing and negative MI beyond 5 dimensions. These failures of kNN are not mere performance degradation but theoretical collapse in high-dimensional regimes. The features selected by PSS also led to higher downstream classification accuracy, confirming that theoretical soundness translates to practical utility.
PSS has limitations common to grid-based methods. The number of partitions grows as $\ell^{d}$, making the method impractical when $d\gg 50$ and samples become too sparse. Our experiments show reliable performance up to $d=40$ with moderate sample sizes, which covers many practical applications. Additionally, as noted in the correlation experiments, extreme dependence (e.g., $\rho\to 1$) leading to manifold collapse remains a challenge for full-dimensional density estimators. From a theoretical perspective, we proved consistency but did not derive convergence rates—characterizing the optimal $\ell(n)$ and bias-variance tradeoff remains open. Other directions include adaptive partition size strategies, extension to mixed discrete-continuous variables, and improved scaling to higher dimensions.
Overall, PSS provides a simple, consistent, and training-free approach to multivariate entropy estimation. Its combination of theoretical guarantees, computational efficiency, and robustness to non-Gaussian distributions makes it useful for applications in feature selection, causal discovery, and representation learning where accurate entropy estimates are needed.
Appendix A Proof of Lemma 1